### v3

# Practical limitations for real-life application of data fission and data thinning in post-clustering differential analysis

Abstract
Post-clustering inference in single-cell RNA sequencing (scRNA-seq) analysis presents significant challenges in controlling Type I error during differential expression analysis. Data fission, a promising approach that aims to split data into two independent parts, relies on strong parametric assumptions of non-mixture distributions that are inherently violated in clustered data. To address this limitation, we introduce conditional data fission, an extension designed to decompose each mixture component into two independent parts. However, we demonstrate that applying such conditional data fission to mixture distributions requires prior knowledge of the clustering structure to ensure valid post-clustering inference. This arises from the need to accurately estimate component-specific scale parameters, which are critical for performing decomposition while maintaining independence. We theoretically quantify how biases in estimating these parameters lead to inflated Type I error rates due to deviations from independence. Given that mixture components are typically unknown in practice, our results underscore the fundamental difficulty of applying data fission in real-world settings, despite its prior proposal as a solution for post-clustering inference.
Keywords: Double-dipping, Mixture model, Post-clustering inference, Type I error, Unsupervised learning
1 Introduction
Clustering encompasses all unsupervised statistical methods that group observations into homogeneous and separated clusters. Widely used in various application fields such as biology or genomics (Jaeger & Banks 2023), clustering plays a major role in revealing or summarizing the signals contained in different types of multivariate data. In single-cell RNA-seq (scRNA-seq) analysis, a high-throughput technique for measuring gene expression at the single-cell level, clustering is typically the first step used to explore cellular heterogeneity and functional diversity (Amezquita et al. 2020). This clustering groups cells according to their gene expression. Afterwards, the differential expression analysis (comparing gene abundance between groups) allows to identify and annotate cellular sub-populations. This leads to the identification of marker genes that could potentially serve as cell-type marker genes (Pullin & McCarthy 2024).
However, such a two-step pipeline for post-clustering differential analysis requires using the same data twice: first to estimate the clusters, and then again to estimate the differences between them and perform significance testing – a procedure sometimes referred to as “double-dipping” (Kriegeskorte et al. 2009). In post-clustering differential analysis, it has been shown that the primary risk of double-dipping is to compromise the control of the Type I error rate of otherwise well-calibrated testing procedures for traditional differential analysis, leading to false positives (Zhang et al. 2019). Fundamentally, uncertainties stemming from cluster analysis results, particularly related to the determination of the optimal number of clusters, could create artificial differences between homogeneous groups of observations. Failure to account for the double use of data during the analysis may lead to erroneously identify those artificial differences as significant (Hivert et al. 2024). Although challenges related to double-dipping are increasingly studied, they remain unresolved in the context of scRNA-seq data analysis (Lähnemann et al. 2020).
To address the double-dipping issue in post-clustering inference, various methodological developments have leveraged the selective inference framework (Fithian et al. 2017), which entails conditioning on the clustering event when deriving test statistics and associated p-values (Gao et al. 2022). For a recent and comprehensive review of existing methods, see Enjalbert Courrech et al. (2025). As an alternative to the selective inference paradigm, Leiner et al. (2023) have drawn inspiration from data splitting (which effectively addresses overfitting issues in machine learning) to break free from the selective inference framework. They propose a method called “data fission”, where the information within each individual observation is split into two independent parts. In theory, the first part can be used for clustering, and then for labeling observations in the second part simply by identifying them with their first matching counterpart. Differential analysis can then be conducted on the remaining information (i.e. the second part). However, in the context of post-clustering differential analysis, it is imperative that the two parts be independent, as each analysis (cluster analysis and differential analysis) must be performed independently to effectively prevent double-dipping and the associated inflation of Type I error. For this, data fission requires strong parametric distributional assumptions with only the Gaussian and Poisson (Neufeld, Gao, Popp, Battle & Witten 2024) distributions ensuring independence between the two fissioned parts. Expanding on this concept, Neufeld, Dharamshi, Gao & Witten (2024) have generalized the fission process by introducing “data thinning”. Building on the same foundational idea, they not only succeed in developing a process capable of decomposing data into more than two parts, but also broaden the spectrum of distributions where independence between each part is provided. This includes the negative binomial distribution, widely used for modeling RNA-seq data.
Data fission and data thinning have emerged as promising alternatives to selective inference for post-clustering inference due to their compatibility with various clustering methods and differential analysis tests. However, they present some inherent limitations that make them difficult to apply for practical post-clustering inference. First, these methods lack results and justification when applied to mixture distributions, which are commonly used to model data with a clustered structure (Macqueen 1967). Consequently, the absence of such results inherently assumes a global null hypothesis of no clusters in the data when applying data fission or data thinning. Additionally, these methods assume prior knowledge of parameters (e.g., variance for the Gaussian distribution or overdispersion for the negative binomial distribution). Although robust estimators for these parameters could theoretically ensure the validity of the method, this further adds complexity for clustered data where each cluster has a different parameter value. In the absence of knowledge about the data structure, specifically potential clusters, the only justifiable estimator is the full-sample one (i.e. computed using all the observations regardless of their mixture component) that fails to correctly estimate the intra-component parameter value.
In this article, we demonstrate that these approaches are not practical for real-world applications. Performing cluster analysis inherently assumes that the data originate from mixture models, contradicting the parametric assumptions of data fission and data thinning, which cannot decompose such distributions. Although one could apply data fission without explicitly accounting for mixture distributions (i.e., marginally to the class membership), we show that this approach leads to incorrect inference due to induced correlations between the two fissioned parts. These correlations arise because marginal data fission relies on an estimate of the overall scale parameter, which is biased relative to the intra-component (conditional) scale parameter. Specifically, we establish a direct link between bias in variance estimation within a Gaussian mixture model and inflation of the Type I error rate in the two-sample t-test (Welch 1947), highlighting the necessity of a robust variance estimator. To ensure independence and valid inference while accounting for mixture structures, the ideal strategy would be to perform decomposition at the component level, that is conditionally on class membership. However, in real-world post-clustering inference, this approach is infeasible as class memberships are unknown and must be estimated through clustering itself, leading to a circular problem.
2 Methods
In the following, capital letters represent random variables with a probability distribution function denoted as $f$, $x_{i}$ denotes a set of $n$ realizations of $X$, and matrices and vectors are in bold.
2.1 Data fission and data thinning
Data fission and data thinning are methods designed to decompose a random variable into two new independent random variables that can ensure the post-clustering inference validity.
Let $\boldsymbol{X}$ be a random variable with a known distribution. Both data fission (Leiner et al. 2023) and its generalization, data thinning (Neufeld, Dharamshi, Gao & Witten 2024), aim to decompose the random variable $\boldsymbol{X}$ into two (or more in data thinning) new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. These new variables are designed to i) retain information from the original variable $\boldsymbol{X}$, and ii) be independent. Of note, data fission can generate pairs of $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ that are not independent. Here, we focus only on the independent cases. The balance between the amount of information from $\boldsymbol{X}$ kept in either $\boldsymbol{X}^{(1)}$ or $\boldsymbol{X}^{(2)}$ is tuned by a hyperparameter $\tau$. Such a decomposition can be performed for various probability distributions of $\boldsymbol{X}$. For data fission, Leiner et al. (2023) identified only two distributions – Gaussian and Poisson – that satisfy the independence requirement. In contrast, Neufeld, Dharamshi, Gao & Witten (2024) established a comprehensive decomposition that ensures independence for all convolution-closed distributions. This encompasses Gaussian, Poisson (Neufeld, Gao, Popp, Battle & Witten 2024) and negative binomial (Neufeld et al. 2023) distributions.
Distribution of $\boldsymbol{X}$ $\boldsymbol{\tau}$ Data fission Data thinning Scale parameter
$\mathcal{P}(\lambda)$ $\tau\in[0,1]$
$W\sim\operatorname{Binom}(X,\tau)$
$X^{(1)}=W$
$X^{(2)}=X-W$
$X^{(1)}\,|\,X=x\sim\operatorname{Binom}(x,\tau)$
$X^{(2)}=X-X^{(1)}$
$\mathcal{N}(\boldsymbol{\mu}_{p},\boldsymbol{\Sigma}_{p\times p})$
$\tau\in]0,+\infty)$
$\tau_{2}\in]0,1[$
$\boldsymbol{W}\sim\mathcal{N}(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{p\times p})$
$\boldsymbol{X}^{(1)}=\boldsymbol{X}+\tau\boldsymbol{W}$
$\boldsymbol{X}^{(2)}=\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{W}$
$\boldsymbol{X}^{(1)}\,|\,\boldsymbol{X}=\boldsymbol{x}\sim\mathcal{N}(\tau_{2}\boldsymbol{x},\tau_{2}(1-\tau_{2})\boldsymbol{\Sigma}_{p\times p})$
$\boldsymbol{X}^{(2)}=\boldsymbol{X}-\boldsymbol{X}^{(1)}$
$\boldsymbol{\Sigma}$
$\operatorname{NegBin}(\mu,\theta)$ $\tau\in[0,1]$ No fission
$X^{(1)}\,|\,X=x\sim\text{BetaBin}(x,\tau\theta,(1-\tau)\theta)$
$X^{(2)}=X-X^{(1)}$
$\theta$
Table 1: Data fission and data thinning decompositions for three usual distributions: Poisson, Gaussian and negative binomial.
The decompositions for the three distributions into two independent random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, are detailed in Table 1. For proofs of independence, please refer to Section $1$ of the Supplementary Materials for the Gaussian case, and to Neufeld, Gao, Popp, Battle & Witten (2024) and Neufeld et al. (2023) for the Poisson and negative binomial distribution, respectively. Data fission or thinning of the Gaussian and negative binomial distributions both require knowledge of a scale parameter (namely $\boldsymbol{\Sigma}$ or $\theta$, respectively) for practical feasibility, as presented in Table 1. Theoretical guarantees, and especially the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$, rely on using the true values of these parameters. Yet, in real-life applications these are unknown and need to be estimated, most likely from the data as described in Neufeld et al. (2023).
2.2 Data fission and thinning for mixture models
2.2.1 Mixture models
Neufeld, Gao, Popp, Battle & Witten (2024) and Neufeld et al. (2023) have introduced data thinning as a framework to address post-clustering inference. However, existing decomposition approaches are limited to probabilistic models that assume homogeneity and fail to account for potential heterogeneity arising from distinct clusters. In other words, these methods rely on a global null hypothesis of no cluster in the data. When true clusters exist, a more suitable modeling approach involves mixture models (Bouveyron et al. 2019), where each component of the mixture corresponds to a distinct cluster.
Finite mixture models provide a probabilistic framework for capturing heterogeneity in clustered data by assuming that each observation arises from one of several latent subpopulations. Formally, let $X_{1},\dots,X_{n}$ be independent observations, and introduce a latent categorical variable $Z_{i}\in\{1,\dots,G\}$ that indicates the class membership of the observation $X_{i}$. The probability that a given observation $X_{i}$ belongs to the cluster $g$ is given by the mixing proportions $\pi_{g}=\mathbb{P}(Z_{i}=g),\text{ such that }\sum_{g=1}^{G}\pi_{g}=1$. Given that $X_{i}$ belongs to the cluster $g$ (i.e. $Z_{i}=g$), its conditional distribution $\mathbb{P}(X_{i}\mid Z_{i}=g,\theta_{g})$ follows a cluster-specific density function $f_{g}(X_{i}\mid\theta_{g})$ , parameterized by $\theta_{g}$.
The marginal distribution of $X_{i}$, obtained by summing over all possible latent classes, is then: $f(X_{i}\mid\theta)=\sum_{g=1}^{G}\pi_{g}f_{g}(X_{i}\mid\theta_{g})$. Mixture models thus offer a principled probabilistic framework for modeling structured heterogeneity, enabling the derivation of data fission and data thinning decompositions in the presence of clustered data.
We will restrict our analysis to the Gaussian setting where the conditional (i.e. cluster-specific) density function is given by:
$\mathbb{P}\left(\boldsymbol{X}_{i}\mid Z_{i}=g,\boldsymbol{\theta}_{g}\right)=f_{g}(\boldsymbol{X}_{i}\mid\boldsymbol{\theta}_{g})=\frac{1}{(2\pi)^{p/2}|\boldsymbol{\Sigma}_{g}|^{1/2}}\exp\left(-\frac{1}{2}(\mathbf{X}_{i}-\boldsymbol{\mu}_{g})^{T}\boldsymbol{\Sigma}_{g}^{-1}(\mathbf{X}_{i}-\boldsymbol{\mu}_{g})\right)$
for $g=1,\dots,G$, where $G>1$ denotes the number of mixture components. Each component is parametrized by its own parameter $\boldsymbol{\theta}_{g}=\left(\boldsymbol{\mu}_{g},\boldsymbol{\Sigma}_{g}\right)$ where $\boldsymbol{\mu}_{g}=\mathbb{E}\left[\boldsymbol{X}_{i}\mid Z_{i}=g\right]\in\mathbb{R}^{p}$ denotes the conditional (i.e. cluster-specific) mean vector and $\boldsymbol{\Sigma}_{g}=\operatorname{\mathbb{V}ar}\left(\boldsymbol{X}_{i}\mid Z_{i}=g\right)\in\mathbb{R}^{p\times p}$ is the conditional (i.e. cluster-specific) variance. Much of the results apply to the negative binomial case, with the overdispersion parameter being analogous to the variance parameter in the Gaussian case.
2.2.2 Conditional (per-component) data fission vs. marginal (global) data fission for mixture models
Since conditionally on $Z_{i}=g$, the random variables $\boldsymbol{X}_{i}$ follow a Gaussian distribution, i.e. $\boldsymbol{X}_{i}\mid Z_{i}=g\sim\mathcal{N}\left(\boldsymbol{\mu}_{g},\boldsymbol{\Sigma}_{g}\right)$, it is possible to conditionally decompose mixture models using traditional data fission or data thinning. Let $\boldsymbol{W}_{1},\dots,\boldsymbol{W}_{n}$ be independent Gaussian random variables such that, for all $g=1,\dots,G$, $\left(\boldsymbol{W}_{i}\mid Z_{i}=g\right)\overset{\text{iid}}{\sim}\mathcal{N}\left(\boldsymbol{0}_{p},\boldsymbol{\Sigma}_{g}\right)$. The data fission process, as described in Table 1, yields the following decomposition of each component of the mixture:
$\left(\boldsymbol{X}^{(1)}_{i}\mid Z_{i}=g\right)=\Big{(}\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i}\mid Z_{i}=g\Big{)}\,\text{and}\,\left(\boldsymbol{X}^{(2)}_{i}\mid Z_{i}=g\right)=\Big{(}\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i}\mid Z_{i}=g\Big{)}.$
This fission decomposition ensures that, for all $g=1,\dots,G$, the random variables $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ are conditionally independent given $Z_{i}$: $\left(\boldsymbol{X}_{i}^{(1)}\perp\!\!\!\perp\boldsymbol{X}^{(2)}_{i}\right)\mid Z_{i}=g$.
Thus, each mixture component is decomposed into two independent Gaussian random variables. However, this conditional data fission does not guarantee the marginal independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$. In particular, we can show that (see Section $2.1$ of Supplementary Materials):
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)=\sum\limits_{g=1}^{G}\pi_{g}\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)^{T}$ (1)
Notably, this conditional independence property of the decomposition holds only when the conditional variances $\boldsymbol{\Sigma}_{g}$ are used to sample the $\boldsymbol{W}_{i}$ and perform the conditional decomposition. However, in the context of post-clustering inference, the class memberships $Z_{i}$ are unknown and must be estimated through clustering. Consequently, the true conditional variances $\boldsymbol{\Sigma}_{g}$ are inaccessible and obtaining reliable estimators for them is a challenge. The only available estimate, without any prior knowledge of $Z_{i}$, is the empirical variance $\widehat{\boldsymbol{\Sigma}}$ which serves as an estimate of the marginal variance $\boldsymbol{\Sigma}$. However, $\widehat{\boldsymbol{\Sigma}}$ is a biased estimate of $\boldsymbol{\Sigma}_{g}.$ It becomes impossible to guarantee the conditional independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}.$ Nevertheless, it is always possible to perform a marginal data fission using the marginal variance $\boldsymbol{\Sigma}$ (or it estimate):
$\boldsymbol{X}^{(1)}_{i}=\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i}\quad\text{and}\quad\boldsymbol{X}^{(2)}_{i}=\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i}\quad\text{with}\quad\boldsymbol{W}_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\boldsymbol{0}_{p},\boldsymbol{\Sigma}\right).$
Such a marginal fission ensures that $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ are uncorrelated but at the cost of introducing conditional and hidden dependencies given by (see Section $2.2$ of Supplementary Materials):
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g\right)=\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma}.$ (2)
Covariance results on conditional and marginal data fission are summarized in Supplementary Table 1.
3 Results
3.1 Limits in practical application of data fission and data thinning for mixture model
In the context of post-clustering inference on mixture models, data fission must be handled with care. The most straightforward approach is marginal data fission, which does not require any knowledge of the latent class memberships or mixture parameters. It relies solely on an estimate of the marginal covariance matrix, $\boldsymbol{\Sigma}$, typically approximated using empirical estimates. However, this simplicity comes at a cost: the two resulting random variables, $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$, are marginally uncorrelated as illustrated in Supplementary Table 1. This implies that the clustering structure present in the original data can only be preserved in one of the two new variables – never both – leading to a partial loss of information across the decomposition. Otherwise, if they both contain information of the latent clustering structure, they would be correlated.
Figure 1: Performance of marginal and conditional data fission in ideal and adverse clustering scenario over $1,000$ simulations under the Gaussian mixture setting. Panel A: Illustration of an ideal configuration where all parameters ($G$, $\pi_{g}$, $\boldsymbol{\mu}_{g}$, $\boldsymbol{\Sigma}_{g}$) are known. Panel B: Empirical statistical power (at the $5\%$ level) of the two-sample $t$-test between Cluster 11 and Cluster 22, as a function of $\tau$ and sample size. Panel C: Adjusted Rand Index (ARI) between $k$-means clustering and true labels after marginal or conditional data fission. Panel D: Illustration of an adverse scenario where a true component is erroneously split into two clusters. Panel E: Empirical Type I error rate (at the $5\%$ level) of the two-sample t-test under spurious Clusters 1 and 2, as a function of $\tau$ and sample size.
This limitation has practical consequences. Even in an ideal scenario where the true mixture parameters are known (Figure 1A), marginal data fission leads to poor transfer of clustering information from $\boldsymbol{X}_{i}^{(1)}$ to $\boldsymbol{X}_{i}^{(2)}$. As a result, downstream inference based on $\boldsymbol{X}_{i}^{(2)}$ suffers from a substantial loss in statistical power (Figure 1B), despite the clustering on $\boldsymbol{X}_{i}^{(1)}$ being nearly perfect (Figure 1C). The inability of $\boldsymbol{X}_{i}^{(2)}$ to benefit from the cluster structure learned in $\boldsymbol{X}_{i}^{(1)}$ stems directly from the marginal uncorrelation between the two fissioned random variables.
In more challenging scenarios where mixture parameters are unknown, marginal fission can even lead to misleading results. Figure 1D illustrates a case in which a single mixture component is wrongly split into two clusters due to imperfect clustering. Under marginal fission, the conditional covariance from equation (2) propagate this incorrect partition – and the artificial differences it introduces – into $\boldsymbol{X}_{i}^{(2)}$. As a result, false positives may arise when testing for differences between clusters that are, in truth, identical. This breakdown of Type I error control, shown in Figure 1E, stems from marginal fission’s conditional dependency property, allowing artifacts in $\boldsymbol{X}_{i}^{(1)}$ to contaminate inference in $\boldsymbol{X}_{i}^{(2)}$.
Conditional data fission offers a more principled alternative. Rather than ignoring the latent structure, it constructs two new random variables that are conditionally independent given the mixture component, while preserving the marginal between-cluster differences between the fissioned random variables. This ensures that clustering performed on $\boldsymbol{X}_{i}^{(1)}$ remains valid for inference in $\boldsymbol{X}_{i}^{(2)}$, without inducing spurious within-cluster correlations. In simulations, conditional fission consistently maintains higher clustering performances and statistical power (Figure 1B) and robust Type I error control (Figure 1E) when clustering is imperfect.
Nevertheless, conditional data fission is not without drawbacks. Its implementation requires knowledge of the component-specific covariance matrices and latent class memberships – quantities that are typically unknown in post-clustering contexts. In our simulations (Figure 1), conditional fission is applied using the true class memberships, providing a best-case benchmark. In real-world applications, however, this latent information must be first estimated through clustering, reintroducing the same post-clustering inference challenges that data fission aims to mitigate.
These simulations also highlighted the strong impact of $\tau$, which controls the proportion of information about $\boldsymbol{X}_{i}$ that is preserved in each fissioned datasets, for both the marginal and conditional data fission. When $\tau$ is small, most of the information is retained in $\boldsymbol{X}_{i}^{(1)}$, facilitating accurate clustering and yielding high Adjusted Rand Index values (Figure 1C). However, this comes at the cost of leaving little information in $\boldsymbol{X}_{i}^{(2)}$, especially under marginal fission, where clustering-related signal is not preserved. This mismatch leads to a sharp drop in statistical power when testing for differences between inferred clusters. Conversely, when $\tau$ is large, more signal is injected into $\boldsymbol{X}_{i}^{(2)}$, improving the potential for downstream testing, but weakening the clustering quality in $\boldsymbol{X}_{i}^{(1)}$. As a result, the clusters compared during testing may not reflect the true latent classes, again reducing power. Therefore, under marginal fission, power is maximized only at intermediate values of $\tau$, where the trade-off between clustering accuracy and test informativeness is partially balanced – but never fully resolved due to the structural disconnect between fissioned random variables. Benefits of conditional data fission are hence particularly pronounced in low-$\tau$ regimes, where signal in $\boldsymbol{X}^{(2)}$ is weak but the preservation of the clustering structure is strengthened by the marginal covariance.
3.2 Type I error relation with bias in the variance estimation
The conditional or marginal independence properties between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ reported in Supplementary Table 1 hold strictly when the true variance matrix–either $\boldsymbol{\Sigma}$ or the component-specific $\boldsymbol{\Sigma}_{g}$–is used for the decomposition. In practice, these quantities are typically unknown and must be estimated from data.
Proposition 1
[…]
Proposition 1 (see also Section 3 of the Supplementary Materials for a proof) shows that replacing the true covariance with an estimate in the data fission process can induce dependencies between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$. A similar result is established by Proposition 10 in Neufeld, Dharamshi, Gao & Witten (2024) for data thinning. In particular, if the estimate $\widehat{\boldsymbol{\Sigma}}$ is biased, it can create spurious correlations between the fissioned datasets, enabling artificial clustering structures–such as the erroneous splitting of a single mixture component into multiple clusters in $\boldsymbol{X}^{(1)}$–to carry over into $\boldsymbol{X}^{(2)}$. As a consequence, Type I error may be inflated during inference, even though testing is conducted exclusively on $\boldsymbol{X}^{(2)}$.
To further elucidate the consequences of variance misestimation in data fission, we derived an analytical expression for the Type I error rate of the $t$-test as a function of the estimation bias. This provides theoretical insight into how variance estimation errors propagate through the data fission process and impact downstream statistical inference. Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it contains no real clusters, and therefore, no actual difference in means exists between subgroups of observations that is not a consequence of the clustering. Let $W_{1},\dots,W_{n}\overset{\text{iid}}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in,]0,+\infty[$. Our aim is to perform data fission of each $X_{i}$, using the $X^{(1)}_{i}$ for $k$-means clustering (with $K=2$) and the $X^{(2)}_{i}$ for differential testing between the two inferred clusters. Here, $b^{2}$ represents any value used as a plug-in for the variance of the $X_{i}$, and in particular, $b^{2}$ may be an estimate obtained for $\sigma^{2}$. Given the generation process of the $X_{i}$, it is established that regardless of the clustering on the $X^{(1)}_{i}$, there should be no mean difference between the estimated clusters on $X^{(2)}_{i}$ as long as independence is achieved. Let $C_{1}$ and $C_{2}$ be the two estimated clusters on the $X^{(1)}_{i}$ with the same intra-cluster variance, which is a reasonable hypothesis with $k$-means clustering as explained in Section $4$ of the Supplementary Materials. Since we are under the null hypothesis of no mean difference between the clusters, the $T$ statistic for the $t$-test between $C_{1}$ and $C_{2}$ using the $X^{(2)}_{i}$ is given by:
$\frac{\overline{X^{(2)}_{C_{1}}}-\overline{X^{(2)}_{C_{2}}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}}\qquad\mbox{where}\qquad\overline{X^{(2)}_{C_{j}}}=\frac{1}{\left|C_{k}\right|}\sum\limits_{i\in C_{k}}X_{i}^{(2)}$ (3)
Here, $s^{2}\left(X^{(2)}\right)$ is their shared intra-cluster variance computed using the $X^{(2)}_{i}$. It can be demonstrated that:
$T\overset{\mathcal{L}}{\sim}\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$
where $\rho=\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=\frac{\left(\sigma^{2}-b^{2}\right)}{\sqrt{\left(\sigma^{2}+\tau^{2}b^{2}\right)\times\left(\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)}}$. Section $4$ of the Supplementary Materials gives details on the derivation of this test statistic and its distribution. The associated Type I error rate for this test is given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of a standard Gaussian distribution $\mathcal{N}(0,1)$.
We validated this theoretical result through numerical simulations, and conducted a detailed exploration of the influence of variance values and sample size on the resulting Type I error rate. We generated $n$ realizations of a Gaussian random variable with a mean $\mu=0$ and variances $\sigma^{2}$, with 1,0001,\!000 Monte Carlo repetitions. Subsequently, we used data fission with varying values of $b^{2}=\widehat{\sigma^{2}}$, obtaining $X^{(1)}$ for $k$-means clustering with $K=2$ and $X^{(2)}$ for testing mean differences between the two estimated clusters. Initially, we examine the impact of the original true variance $\sigma^{2}$ by considering $\sigma^{2}$ values of $\{0.01,0.25,1,4\}$ for a fixed sample size of $n=100$. Figure 2A, illustrates the relationship between the bias in estimating $\sigma^{2}$ and the Type I error rate, demonstrating a consistent agreement with the derived theoretical error rate.We further documented the behavior of the Type I error for a fixed $\sigma^{2}=1$ with varying sample sizes $n\in\{50,100,200,100,500,1,\!000\}$ in Figure 2B, which shows the expected impact of the sample size on the Type I error rate. These results show the critical importance of accurately estimating the variance to achieve a well-calibrated Type I error rate with data fission.
Figure 2: Impact of variance estimation on Type I error in data fission. Type I error as a function of the relative bias colored by (A) true data variance, or (B) sample size.
This issue with variance estimation fully accounts for the conditional correlations observed under marginal data fission. Given class membership, the Gaussian assumption is satisfied, and the correct covariance structure is component-specific. However, marginal fission relies on a global covariance estimate, $\boldsymbol{\Sigma}$, which averages over the component-specific covariances $\boldsymbol{\Sigma}_{g}$, $g=1,\dots,G$. This averaging introduces bias and leads to a misspecified covariance structure. As a result, dependencies between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ emerge at the component level, thereby compromising inference. By contrast, conditional data fission–when implemented with accurate, component-specific variance estimates–preserves the intended conditional independence structure and is therefore better suited for mixture settings.
3.3 Application to single-cell RNA-seq data analysis
Single-cell RNA-seq (scRNA-seq) data analysis pipelines often involve an initial cluster analysis followed by differential analysis to identify marker genes and annotate cell populations based on gene expression. Given their overdispersed count nature, the negative binomial distribution is favored over the Gaussian distribution to model scRNA-seq data. Unfortunately, the negative binomial distribution also raises challenges for data thinning. The overdispersion parameter $\theta$ plays a role similar to the variance parameter in the Gaussian distribution, as it is considered known. Therefore, the quality of its estimation directly impacts $\operatorname{Cov}\left(X^{(1)},X^{(2)}\right)=\tau(1-\tau)\frac{\mu^{2}}{\theta}\left(1-\frac{\theta+1}{\widehat{\theta}+1}\right)$, where $\widehat{\theta}$ is an estimate of $\theta$ (Neufeld, Dharamshi, Gao & Witten 2024). Moreover, in negative binomial mixture models, conditional data thinning is also preferable to marginal thinning due to the component-specific nature of the overdispersion parameter (Li et al. 2018). However, applying conditional thinning requires accurate estimates of these parameters, which is challenging in practice since class memberships are unknown and must be inferred. This makes it difficult to ensure conditional independence and control Type I error in post-thinning analyses.
We further illustrate the necessity of applying conditional data thinning (with the associated conditional overdispersion) to ensure Type I error control in the post-clustering inference setting with numerical simulations. We generated $n=100$ observations from a two-component negative binomial mixture: $0.5\text{NegBin}(\mu_{1},\theta_{1})+0.5\text{NegBin}(\mu_{2},\theta_{2})$ with component parameters $(\mu_{1},\theta_{1})=(5,5)$ and $(\mu_{2},\theta_{2})=(60,40)$. We conducted similar post-clustering inference as in Figure 1. In Figure 3A, applying the $k$-means with $G=3$ clusters reveals that the first mixture component is erroneously split into 2 clusters (labeled $C_{1}$ and $C_{3}$). We then assessed the Type I error rate associated with the Wilcoxon test between these spurious clusters when applying both marginal and conditional data thinning. First, we applied conditional data thinning using oracle estimates $\widetilde{\theta}_{g}$, $g=1,2$, representing true intra-component overdispersion parameters (infeasible in real-life application where the true class memberships are unknown). We compared those results with marginal data thinning with its associated marginal $\widehat{\theta}$. Of note, all the overdispersion estimations were performed using Maximum Likelihood Estimation. Figure 3B presents the QQ-plot against the Uniform distribution of the associated Wilcoxon p-values across 1,0001,\!000 simulations. Similarly to the Gaussian setting, achieving uniformly distributed p-values is only possible through conditional data thinning using oracle conditional overdispersion estimates $\widetilde{\theta}_{g}$. Marginal negative binomial data thinning does not guarantee conditional independence between $X^{(1)}$ and $X^{(2)}$, leading to inflated Type I error rates in post-clustering inference – particularly when a true component is erroneously split into multiple clusters.
Figure 3: Challenges in estimating overdispersion, a gene-specific parameter, for negative binomial data thinning. A: spurious clustering in simulated data. B: $Q$-plot of Wilcoxon p-values when testing between two spurious clusters. C: concordance between gene-wise overdispersions estimated within 4 different cell populations and in neutrophils.
Extending our investigation from simulated scenarios to real-life applications, we used a single-cell RNA-seq dataset from the Tabula Sapiens Consortium (The Tabula Sapiens Consortium* 2022) to explore the practical challenges of estimating overdispersion. Our analysis focused on five distinct cell populations from bone marrow: 2,5602,\!560 neutrophils, $105$ macrophages, $386$ monocytes, $454$ granulocytes, and $833$ CD4 T cells – all collected from a single donor (donor TSP14). In this controlled setting, where cell types were known, we succeeded in estimating the overdispersion of 8,3338,\!333 genes for each cell type using the variance stabilizing transformation implemented in the sctransform package (Choudhary & Satija 2022). Figure 3C illustrates the comparison of gene overdispersion when estimated solely in neutrophils versus the overdispersion estimated for the same genes in the other four cellular populations. Root Mean Squared Error (RMSE) values were computed to quantify the agreement between estimations. Our findings reveal that overdispersion is specific to each cell population, as evidenced by a notable deviation from the diagonal and relatively high RMSE values. This highlights the need for conditional data thinning in such an application. Combined with our simulation studies, these results demonstrate how Type I error can easily be inflated in real-life applications of data thinning for scRNA-seq data analysis, particularly due to the difficulty in providing an unbiased estimate of gene overdispersion to be used in conditional data thinning.
4 Discussion
We highlight here the practical limitations inherent in data fission and its extension, data thinning, for post-clustering inference challenges. An important issue is the assumption of a homogeneous data distribution, which implies the absence of true clusters in the data. To address this limitation and adapt to scenarios with true classes, a shift towards mixture models becomes imperative. However, these models lack a predefined decomposition through data fission or thinning. We have shown that applying data fission or thinning without accounting for the mixture structure – that is, using marginal approaches – leads to inflated Type I error rates and quite low statistical power in post-clustering inference.
We have proposed conditional data fission and data thinning, which perform the decomposition at the component level, and demonstrated their theoretical validity. However, theses approaches rely on prior knowledge of the mixture components scale parameters, such as variances in the Gaussian distribution or the overdispersion in the negative binomial distribution. However in real-life applications, these parameters are unknown. Adequately estimating these parameters becomes intricate in the presence of true clusters, given their component-specific nature; meanwhile the quality of the estimation of those parameters is directly linked with the covariance between the new random variables, $X^{(1)}$ and $X^{(2)}$, decomposing the original data. Only unbiased estimation of these parameters ensures the conditional independence between $X^{(1)}$ and $X^{(2)}$. That independence is paramount for post-clustering inference to adequately control the Type I error rate. Thus, although conditional data fission offers a promising solution for valid post-clustering inference, it creates a circular dependency. The clusters must be known in advance to perform the decomposition, even though this decomposition is intended to enable valid inference about such estimated clusters (see Supplementary Figure 1).
In the Gaussian framework, we theoretically quantify the relationship between the relative bias in variance estimation and the associated Type I error in post-clustering $t$-tests. In practice, our simulation results suggest that a small relative bias can be acceptable while still achieving effective Type I error control. These first results pave the way for defining a principled approach to tuning the hyperparmater $\tau$ in data fission and data thinning for post-clustering inference to optimize statistical power.
Finally, we demonstrate that the results derived from the Gaussian distribution context are readily applicable to the negative binomial distribution commonly employed in modeling RNA-seq data. Specifically, we illustrate on real data that overdispersion is also component-specific. Therefore, without knowing the true class memberships, data thinning cannot ensure the needed independence between clustering and differential testing.
To avoid the need for prior knowledge of component-specific variances in Gaussian mixture settings, one may consider adopting a heteroscedastic model with observation-level means and variances. Within this framework, variances can be approximated using plug-in estimators, such as non-parametric estimates of local variability. This approach is more consistent with the distributional assumptions underlying data fission and thinning. Moreover, such heteroscedastic models naturally encompass mixture models as a special case, where the local variance reflects the component-specific structure. Thus, they might hold a path for approximating conditional data fission, without requiring explicit knowledge of the mixture parameters. Nevertheless, their practical deployment presents notable challenges. Chief among them is the sensitivity of non-parametric variance estimates to the choice of bandwidth. Ideally, the bandwidth should be small enough to capture only observations from the same component. Yet, this implicitly assumes knowledge of the true mixture structure, and its effectiveness ultimately depends on the degree of separation between components.
In practice, the application of data fission or data thinning for post-clustering inference appears to be akin to a self-referential loop generating circular reasoning. While introduced as a solution for addressing challenges in post-clustering inference, all strategies that could theoretically ensure independence between the two stages of the analysis ultimately rely on knowledge of the true, but unknown, class memberships. Despite its conceptual appeal, the practical utility of these methods for post-clustering inference remains limited to extreme cases with extremely high signal vs. noise ratios, emphasizing the need for alternative methodologies that can navigate the complexities of unknown class structures more effectively.
All codes and data needed to reproduce the results presented here are openly accessible from Zenodo with DOI 10.5281/zenodo.11207776.
5 Acknowledgements
BH is supported partly by the Digital Public Health Graduate’s school, funded by the PIA $3$ (Investments for the Future – Project reference: 17-EURE-0019). The work was supported through the DESTRIER Inria Associate-Team from the Inria@SiliconValley program (analytical code: DRI-012215), and by the CARE project funded from the Innovative Medicines Initiative 2 Joint Undertaking (JU) under grant agreement no IMI2-101005077. The JU receives support from the European Union’s Horizon $2020$ research and innovation programme and EFPIA and Bill &\& Melinda Gates Foundation, Global Health Drug Discovery Institute, University of Dundee. This work has received funding from the European Union’s Horizon $2020$ research and innovation programme under EHVA grant agreement no H2020-681032. This study was carried out in the framework of the University of Bordeaux’s France 2030 program / RRI PHDS. This work benefited from State aid managed by the Agence Nationale de la Recherche under the France 2030 program, reference AI4scMED ANR-22-PESN-0002. Computer time for this study was provided by the computing facilities MCIA (Mésocentre de Calcul Intensif Aquitain) of the Université de Bordeaux and of the Université de Pau et des Pays de l’Adour. The authors disclose the use of generative AI (ChatGPT, OpenAI, 2025) to help refine the phrasing of some sections of the manuscript.
Supplementary Materials
Appendix 1 Independence proof of the Gaussian process
Let $\boldsymbol{X}\sim\mathcal{N}\left(\boldsymbol{\mu},\boldsymbol{\Sigma}_{p\times p}\right)$. Considering the fission process for Gaussian data described in Table $1$, we can decompose $\boldsymbol{X}$ into two new random variables $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$ using a new random variable $\boldsymbol{W}\sim\mathcal{N}\left(\boldsymbol{0},\boldsymbol{\Sigma}\right)$. It follows form this decomposition that:
$\boldsymbol{X}^{(1)}\sim\mathcal{N}\left(\boldsymbol{\mu},\left(1+\tau^{2}\right)\boldsymbol{\Sigma}\right)\quad\text{and}\quad\boldsymbol{X}^{(2)}\sim\mathcal{N}\left(\boldsymbol{\mu},\left(1+\frac{1}{\tau^{2}}\right)\boldsymbol{\Sigma}\right)$
Moreover, we have $\boldsymbol{X}^{(1)}\perp\!\!\!\perp\boldsymbol{X}^{(2)}$. Indeed, we have:
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})$ $=\mathbb{E}\left[\left(\boldsymbol{X}^{(1)}-\mathbb{E}[\boldsymbol{X}^{(1)}]\right)\left(\boldsymbol{X}^{(2)}-\mathbb{E}[\boldsymbol{X}^{(2)}]\right)^{t}\right]$
$=\mathbb{E}\left[\left(\boldsymbol{X}^{(1)}-\boldsymbol{\mu}\right)\left(\boldsymbol{X}^{(2)}-\boldsymbol{\mu}\right)^{t}\right]\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{X}^{(1)}]=\mathbb{E}[\boldsymbol{X}^{(2)}]=\boldsymbol{\mu}$
$=\mathbb{E}\left[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}-\boldsymbol{X}^{(1)}\boldsymbol{\mu}^{t}-\boldsymbol{\mu}{\boldsymbol{X}^{(2)}}^{t}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}\right]$
$=\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]-\mathbb{E}[{\boldsymbol{X}^{(1)}}]\boldsymbol{\mu}^{t}-\boldsymbol{\mu}\mathbb{E}[{\boldsymbol{X}^{(2)}}^{t}]+\boldsymbol{\mu}\boldsymbol{\mu}^{t}$
$=\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]-\boldsymbol{\mu}\boldsymbol{\mu}^{t}-\boldsymbol{\mu}\boldsymbol{\mu}^{t}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}$
$=\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]-\boldsymbol{\mu}\boldsymbol{\mu}^{t}$
Moreover, we have:
$\mathbb{E}[\boldsymbol{X}^{(1)}{\boldsymbol{X}^{(2)}}^{t}]$ $=\mathbb{E}\left[\left(\boldsymbol{X}+\tau\boldsymbol{W}\right)\left(\boldsymbol{X}-\frac{1}{\tau}\boldsymbol{W}\right)^{t}\right]$
$=\mathbb{E}\left[\boldsymbol{X}\boldsymbol{X}^{t}-\frac{1}{\tau}\boldsymbol{X}\boldsymbol{W}^{t}+\tau\boldsymbol{W}\boldsymbol{X}^{t}-\boldsymbol{W}\boldsymbol{W}^{t}\right]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}\boldsymbol{W}^{t}]+\tau\mathbb{E}[\boldsymbol{W}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\frac{1}{\tau}\mathbb{E}[\boldsymbol{X}]\mathbb{E}[\boldsymbol{W}^{t}]+\tau\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]\qquad\mbox{since}\qquad\boldsymbol{X}\perp\!\!\!\perp\boldsymbol{W}$
$=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{W}]=0$
But we also have:
$\mathbb{V}\text{ar}(\boldsymbol{X})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{X}]\mathbb{E}[\boldsymbol{X}^{t}]=\boldsymbol{\Sigma}$
$\iff\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]=\boldsymbol{\Sigma}+\mathbb{E}[\boldsymbol{X}]\mathbb{E}[\boldsymbol{X}^{t}]$
$\iff\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]=\boldsymbol{\Sigma}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{X}]=\boldsymbol{\mu}$
and
$\mathbb{V}\text{ar}(\boldsymbol{W})=\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]-\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}$
$\iff\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}+\mathbb{E}[\boldsymbol{W}]\mathbb{E}[\boldsymbol{W}^{t}]$
$\iff\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]=\boldsymbol{\Sigma}\qquad\mbox{since}\qquad\mathbb{E}[\boldsymbol{W}]=0$
So finally,
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\mathbb{E}[\boldsymbol{X}\boldsymbol{X}^{t}]-\mathbb{E}[\boldsymbol{W}\boldsymbol{W}^{t}]-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=\boldsymbol{\Sigma}+\boldsymbol{\mu}\boldsymbol{\mu}^{t}-\boldsymbol{\Sigma}-\boldsymbol{\mu}\boldsymbol{\mu}^{t}=0$ (4)
Appendix 2 Covariance Structures in Conditional and Marginal Gaussian Data Fission
In the following, we assume that $\boldsymbol{X}_{i}$ is distributed according to a Gaussian mixture with conditional variance $\boldsymbol{\Sigma}_{g}$ and marginal variance $\boldsymbol{\Sigma}$.
2.1 Conditional Data Fission
Conditional data fission guarantees conditional independence between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ due to the properties of Gaussian decomposition (see Section $1$ of Supplementary Materials). However, marginal dependence may still arise, as shown below:
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)$ $=\mathbb{E}\left[\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}\right)\right]+\operatorname{Cov}\left(\mathbb{E}\left[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}\right],\mathbb{E}\left[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}\right]\right)$
$=\operatorname{Cov}\left(\mathbb{E}[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}],\mathbb{E}[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}]\right)\quad\text{(by conditional independence)}$
$=\operatorname{Cov}(\boldsymbol{\mu}_{Z},\boldsymbol{\mu}_{Z})$
$=\mathbb{V}\text{ar}(\boldsymbol{\mu}_{Z}),$
where $\boldsymbol{\mu}_{Z}=\mathbb{E}[\boldsymbol{X}_{i}^{(1)}\mid Z_{i}]=\mathbb{E}[\boldsymbol{X}_{i}^{(2)}\mid Z_{i}]$. Since $Z_{i}$ is discrete, we have $\boldsymbol{\mu}_{Z}=\boldsymbol{\mu}_{g}$ with probability $\pi_{g}$, so:
$\mathbb{E}[\boldsymbol{\mu}_{Z}]=\sum_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g},\quad\text{and}\quad\mathbb{V}\text{ar}(\boldsymbol{\mu}_{Z})=\sum_{g=1}^{G}\pi_{g}(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])^{T}.$
Therefore, the marginal covariance between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$ in conditional fission equals the variance of the component means:
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)=\sum_{g=1}^{G}\pi_{g}(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])(\boldsymbol{\mu}_{g}-\mathbb{E}[\boldsymbol{\mu}_{Z}])^{T}.$
2.2 Marginal Data Fission
Marginal covariance
We now demonstrate that marginal data fission induces no marginal correlations between $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$. Let:
$\boldsymbol{X}_{i}^{(1)}=\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\quad\boldsymbol{X}_{i}^{(2)}=\boldsymbol{X}_{i}-\frac{1}{\tau}\boldsymbol{W}_{i},$
where $\boldsymbol{X}_{i}$ has covariance matrix $\boldsymbol{\Sigma}$ and $\boldsymbol{W}_{i}\overset{\text{iid}}{\sim}\mathcal{N}(\mathbf{0},\boldsymbol{\Sigma})$, with $\boldsymbol{X}_{i}\perp\!\!\!\perp\boldsymbol{W}_{i}$.
$\operatorname{Cov}(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)})$ $=\operatorname{Cov}(\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\boldsymbol{X}_{i}-\tfrac{1}{\tau}\boldsymbol{W}_{i})$
$=\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{X}_{i})+\tau\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{X}_{i})-\tfrac{1}{\tau}\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{W}_{i})-\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{W}_{i})$
$=\boldsymbol{\Sigma}-\boldsymbol{\Sigma}$
$=\mathbf{0}.$
Conditional covariance
Next, we quantify the conditional dependence given the class membership $Z_{i}=g$:
$\operatorname{Cov}(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g)$ $=\operatorname{Cov}(\boldsymbol{X}_{i}+\tau\boldsymbol{W}_{i},\boldsymbol{X}_{i}-\tfrac{1}{\tau}\boldsymbol{W}_{i}\mid Z_{i}=g)$
$=\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{X}_{i}\mid Z_{i}=g)+\tau\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{X}_{i}\mid Z_{i}=g)$
$\quad-\tfrac{1}{\tau}\operatorname{Cov}(\boldsymbol{X}_{i},\boldsymbol{W}_{i}\mid Z_{i}=g)-\operatorname{Cov}(\boldsymbol{W}_{i},\boldsymbol{W}_{i}\mid Z_{i}=g)$
$=\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma},$
where $\boldsymbol{\Sigma}_{g}$ denotes the class-specific covariance. Thus, unless all components share the same covariance structure ($\boldsymbol{\Sigma}_{g}=\boldsymbol{\Sigma}$), marginal data fission introduces conditional correlation despite no marginal dependence.
Appendix 3 Impact of covariance estimation under the independence between $\boldsymbol{X}^{(1)}$ and $\boldsymbol{X}^{(2)}$
[…]
$\text{Cov}(\boldsymbol{X}^{(1)},\boldsymbol{X}^{(2)})=\boldsymbol{\Sigma}-\widehat{\boldsymbol{\Sigma}}$
Appendix 4 Derivation of $t$-test statistic under the null of no-cluster in the univariate data fission post-clustering setting
Let $X_{1},\dots,X_{n}$ be $n$ independent and identically distributed random variables such that, for all $i=1,\dots,n$, $X_{i}\overset{\text{iid}}{\sim}\mathcal{N}\left(\mu,\sigma^{2}\right)$. Here, $n$ represents the sample size. Let $W_{1},\dots,W_{n}\overset{\text{iid}}{\sim}\mathcal{N}\left(0,b^{2}\right)$ and $\tau\in(0,+\infty)$. Here, $b^{2}$ represents any value used as a plug-in for the variance of $X$, and in particular, $b^{2}$ can be an estimate obtained for $\sigma^{2}$. For all $i=1,\dots,n$, the splitting process of $X_{i}$ is given by:
$X^{(1)}_{i}=X_{i}+\tau W_{i}\qquad\text{and}\qquad X^{(2)}_{i}=X_{i}-\frac{1}{\tau}W_{i}$
We can immediately deduce the marginal distributions of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ for all $i=1,\dots,n$, thanks to the independence between $X_{i}$ and $W_{i}$:
$X^{(1)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}+\tau^{2}b^{2}\right)\qquad\text{and}\qquad X^{(2)}_{i}\sim\mathcal{N}\left(\mu,\sigma^{2}+\frac{1}{\tau^{2}}b^{2}\right)$ (5)
We denote $\sigma^{2}_{X^{(1)}}=\sigma^{2}+\tau^{2}b^{2}$ and $\sigma^{2}_{X^{(2)}}=\sigma^{2}+\frac{1}{\tau^{2}}b^{2}$ as the respective variances of $X^{(1)}_{i}$ and $X^{(2)}_{i}$ above.
In the context of data fission to address the challenges of post-clustering inference, a clustering algorithm is applied to the observations of $\boldsymbol{X}^{(1)}$. Without loss of generality, we assume that the clustering algorithm applied to the realizations separates $X^{(1)}_{1},\dots,X^{(1)}_{n}$ into two clusters $C_{1}$ and $C_{2}$ around $\mu$ (which is typically the case with the $k$-means algorithm or with a two-component Gaussian mixture model with homogeneous variance when $n$ is sufficiently large). We also assume that these two clusters have the same size and the same variance.
Thus, the clusters $C_{1}$ and $C_{2}$ can be expressed as:
$C_{1}=\left\{i=1,\dots,n:X^{(1)}_{i}>\mu\right\}\qquad\text{and}\qquad C_{2}=\left\{i=1,\dots,n:X^{(1)}_{i}\leq\mu\right\}$
[…]
$\mathbb{E}\left[X^{(1)}_{i}|X^{(1)}_{i}>\mu\right]=\mu+\sqrt{\frac{2\sigma^{2}_{X^{(1)}}}{\pi}}\quad\text{and}\quad\operatorname{Var}\left(X^{(1)}_{i}|X^{(1)}_{i}>\mu\right)=\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(1)}}$
Since the cluster $C_{2}=\left\{i=1,\dots,n:X^{(1)}_{i}\leq\mu\right\}$ simply represents the cluster on the other side of the mean $\mu$, we similarly have:
$\mathbb{E}\left[X^{(1)}_{i}|X^{(1)}_{i}\leq\mu\right]=\mu-\sqrt{\frac{2\sigma^{2}_{X^{(1)}}}{\pi}}\quad\text{and}\quad\operatorname{Var}\left(X^{(1)}_{i}|X^{(1)}_{i}\leq\mu\right)=\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(1)}}$
In the context of post-clustering inference, hypothesis tests are performed on the other part of the information, contained in this case in $\boldsymbol{X}^{(2)}$. We are interested in performing a two-sample $t$-test to evaluate a potential difference in means on $\boldsymbol{X}^{(2)}$ according to the groups defined by the two clusters $C_{1}=\left\{i=1,\dots,n:X_{i}^{(1)}>\mu\right\}$ and $C_{2}=\left\{i=1,\dots,n:X_{i}^{(1)}\leq\mu\right\}$. Thus, we focus on the following hypotheses:
$\mathcal{H}_{0}:\mu_{C_{1}}=\mu_{C_{2}}\qquad\text{vs}\qquad\mathcal{H}_{1}:\mu_{C_{1}}\neq\mu_{C_{2}}$
where $\mu_{C_{1}}=\mathbb{E}\left[X_{i}^{(2)}|X^{(1)}_{i}>\mu\right]$ and $\mu_{C_{2}}=\mathbb{E}\left[X_{i}^{(2)}|X^{(1)}_{i}\leq\mu\right]$ are the means of $X^{(2)}_{i}$ in cluster $C_{1}$ and cluster $C_{2}$, respectively.
Since we have assumed that the two resulting clusters have equal variances (and the same size $n/2$), we denote the common variance as $s^{2}(X^{(2)})=\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}>\mu\right)=\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}\leq\mu\right)$. The corresponding test statistic is then given by:
$T=\frac{\overline{X^{(2)}_{C_{1}}}-\overline{X^{(2)}_{C_{2}}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}}\quad\text{where}\quad\overline{X^{(2)}_{C_{k}}}=\frac{1}{n/2}\sum\limits_{i\in C_{k}}X^{(2)}_{i}\quad\text{for}\quad k=1,2$
Although each $X^{(2)}_{i}$ is Gaussian, this is no longer true conditionally on the clusters, i.e., on $X^{(1)}_{i}>\mu$ for $C_{1}$ (or on $X^{(1)}_{i}\leq\mu$ for $C_{2}$). However, when $n$ is sufficiently large, we can apply the Central Limit Theorem, which gives us:
$\overline{X^{(2)}_{C_{1}}}-\overline{X^{(2)}_{C_{2}}}\overset{\mathcal{L}}{\sim}\mathcal{N}\left(\mu_{C_{1}}-\mu_{C_{2}},\frac{4}{n}s^{2}\left(X^{(2)}\right)\right)$
The asymptotic distribution of our test statistic $T$ is therefore:
$T=\frac{\overline{X^{(2)}_{C_{1}}}-\overline{X^{(2)}_{C_{2}}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}}\overset{\mathcal{L}}{\sim}\mathcal{N}\left(\frac{\mu_{C_{1}}-\mu_{C_{2}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}},1\right)$ (6)
This asymptotic distribution therefore depends on three quantities: $\mu_{C_{1}}$, $\mu_{C_{2}}$, and $s^{2}(X^{(2)})$, which can be computed. By the law of total expectation, we observe that:
$\mu_{C_{1}}=\mathbb{E}\left[X^{(2)}_{i}|X^{(1)}_{i}>\mu\right]=\mathbb{E}\left[\mathbb{E}\left[X^{(2)}_{i}|X^{(1)}_{i}\right]|X^{(1)}_{i}>\mu\right]$ (7)
Since $X^{(1)}_{i}$ and $X^{(2)}_{i}$ are two Gaussian random variables, for all $i=1,\dots,n$, we have the following bivariate Gaussian vector:
$\begin{pmatrix}X^{(1)}_{i}\\ X^{(2)}_{i}\end{pmatrix}\sim\mathcal{N}_{2}\left(\begin{pmatrix}\mu\\ \mu\end{pmatrix},\begin{pmatrix}\sigma^{2}_{X^{(1)}}&\rho\sigma_{X^{(1)}}\sigma_{X^{(2)}}\\ \rho\sigma_{X^{(1)}}\sigma_{X^{(2)}}&\sigma_{X^{(2)}}^{2}\end{pmatrix}\right)$
where $\rho=\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)$. Using the properties of multivariate Gaussian distributions, we can deduce the conditional distribution of $X^{(2)}_{i}|X^{(1)}_{i}$, which for all $i=1,\dots,n$, is:
[…]
which simplifies to:
$X^{(2)}_{i}|X^{(1)}_{i}\sim\mathcal{N}\left(\mu+\rho\frac{\sigma_{X^{(2)}}}{\sigma_{X^{(1)}}}\left(X^{(1)}_{i}-\mu\right),\sigma^{2}_{X^{(2)}}\left(1-\rho^{2}\right)\right)$ (9)
Substituting the expectation of $X^{(2)}_{i}|X^{(1)}_{i}$ from above into equation (7), we obtain:
$\mu_{C_{1}}$ $=\mathbb{E}\left[\mu+\rho\frac{\sigma_{X^{(2)}}}{\sigma_{X^{(1)}}}\left(X^{(1)}_{i}-\mu\right)|X_{i}^{(1)}>\mu\right]$
$=\mu+\rho\frac{\sigma_{X^{(2)}}}{\sigma_{X^{(1)}}}\left(\mathbb{E}\left[X_{i}^{(1)}|X_{i}^{(1)}>\mu\right]-\mu\right)$
$=\mu+\rho\frac{\sigma_{X^{(2)}}}{\sigma_{X^{(1)}}}\left(\mu+\sqrt{\frac{2\sigma^{2}_{X^{(1)}}}{\pi}}-\mu\right)$
$=\mu+\rho\sqrt{\frac{2}{\pi}\sigma^{2}_{X^{(2)}}}$
By an identical reasoning, we find:
$\mu_{C_{2}}=\mathbb{E}\left[X^{(2)}_{i}|X^{(1)}_{i}\leq\mu\right]=\mu-\rho\sqrt{\frac{2}{\pi}\sigma^{2}_{X^{(2)}}}$
Finally, using the law of total variance, we have:
$\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}>\mu\right)$ $=\mathbb{E}\left[\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}\right)|X_{i}^{(1)}>\mu\right]+\operatorname{Var}\left(\mathbb{E}\left[X^{(2)}_{i}|X^{(1)}_{i}\right]|X_{i}^{(1)}>\mu\right)$
From equation (9), we first observe that:
$\mathbb{E}\left[\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}\right)\right|X_{i}^{(1)}>\mu]$ $=\mathbb{E}\left[\sigma^{2}_{X^{(2)}}\left(1-\rho^{2}\right)|X^{(1)}_{i}>\mu\right]$
$=\sigma^{2}_{X^{(2)}}\left(1-\rho^{2}\right)$
and then that:
$\operatorname{Var}\left(\mathbb{E}\left[X^{(2)}_{i}|X^{(1)}_{i}\right]|X_{i}^{(1)}>\mu\right)$ $=\operatorname{Var}\left(\mu+\rho\frac{\sigma_{X^{(2)}}}{\sigma_{X^{(1)}}}\left(X^{(1)}_{i}-\mu\right)|X^{(1)}_{i}>\mu\right)$
$=\rho^{2}\frac{\sigma^{2}_{X^{(2)}}}{\sigma^{2}_{X^{(1)}}}\operatorname{Var}\left(X_{i}^{(1)}|X^{(1)}_{i}>\mu\right)$
$=\rho^{2}\frac{\sigma^{2}_{X^{(2)}}}{\sigma^{2}_{X^{(1)}}}\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(1)}}$
$=\rho^{2}\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(2)}}$
Finally, we obtain:
$s^{2}\left(X^{(2)}\right)$ $=\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}>\mu\right)$
$=\sigma^{2}_{X^{(2)}}\left(1-\rho^{2}\right)+\rho^{2}\left(1-\frac{2}{\pi}\right)\sigma^{2}_{X^{(2)}}$
$=\sigma^{2}_{X^{(2)}}\left(1-\frac{2}{\pi}\rho^{2}\right)$
By a similar reasoning, we find:
$\operatorname{Var}\left(X^{(2)}_{i}|X^{(1)}_{i}\leq\mu\right)=\sigma^{2}_{X^{(2)}}\left(1-\frac{2}{\pi}\rho^{2}\right),$
which, fortunately, confirms our initial assumption of equal intra-cluster variances. Thus, we can finally compute:
$\mathbb{E}\left[T\right]$ $=\frac{\mu_{C_{1}}-\mu_{C_{2}}}{\sqrt{\frac{4s^{2}\left(X^{(2)}\right)}{n}}}=\frac{\mu+\rho\sqrt{\frac{2}{\pi}\sigma^{2}_{X^{(2)}}}-\left(\mu-\rho\sqrt{\frac{2}{\pi}\sigma^{2}_{X^{(2)}}}\right)}{\sqrt{\frac{4\sigma^{2}_{X^{(2)}}\left(1-\frac{2}{\pi}\rho^{2}\right)}{n}}}=\frac{\rho\sqrt{n}\sqrt{\frac{2}{\pi}}}{\sqrt{1-\frac{2}{\pi}\rho^{2}}},$
and we ultimately find:
$T\overset{\mathcal{L}}{\sim}\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right).$ (10)
Recall that the sample $X_{1},\dots,X_{n}$ is normally distributed: it does not contain any true clusters, and thus no real difference in means exists between subgroups of observations other than what is due to clustering. Therefore, the test statistic $T$ should be under $\mathcal{H}_{0}$, and thus centered around 0. In our result in (10), we observe a deviation of the distribution of $T$ from 0, quantified by $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$. The Type I error at level $\alpha$ associated with this test is then given by $1-F(q_{\alpha/2})+F(-q_{\alpha/2})$, where $F$ is the cumulative distribution function of the normal distribution $\mathcal{N}\left(\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}},1\right)$ and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the standard normal distribution $\mathcal{N}(0,1)$.
Here we assumed that all variances were known, and thus $s^{2}\left(X^{(2)}\right)$ was known as well. Therefore, it was possible to calculate the distribution of the test statistic for the $Z$ test to compare means between two samples. However, this result extends easily to the more practical case where $s^{2}\left(X^{(2)}\right)$ is unknown and an estimate $\widehat{s^{2}}\left(X^{(2)}\right)$ is used instead. In this case, still assuming that variances are equal between the two clusters, the distribution of the test statistic $T$ follows a Student’s $t$ distribution $\mathcal{T}(n-2)$ (due to the uncertainties associated with estimating this common variance). The associated Type I error then becomes: $1-F_{\mathcal{T}}(q_{\alpha/2})+F_{\mathcal{T}}(-q_{\alpha/2})$, where $F_{\mathcal{T}}$ is the cumulative distribution function of the non-central Student’s $t$ distribution with mean $\frac{\rho\sqrt{n}}{\sqrt{\frac{\pi}{2}-\rho^{2}}}$ and n2n-2 degrees of freedom, and $q_{\alpha/2}$ is the quantile of order $\alpha/2$ of the Student’s $t$ distribution with n2n-2 degrees of freedom.
This result underscores the crucial importance of precise variance estimation for the application of data fission. Indeed, for the test to be valid, that is, for the Type I error to be controlled at level $\alpha$, the distribution of the test statistic must be centered at 0. This implies that:
$\rho\sqrt{n}=0$ $\iff\rho=0$
$\iff\operatorname{Cor}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=0\quad\forall i=1,\dots,n$
$\iff\frac{\operatorname{Cov}\left(X^{(1)}_{i},X^{(2)}_{i}\right)}{\sigma_{X^{(1)}}\sigma_{X^{(2)}}}=0\quad\forall i=1,\dots,n$
$\iff\operatorname{Cov}\left(X^{(1)}_{i},X^{(2)}_{i}\right)=0\quad\forall i=1,\dots,n$
$\iff\sigma^{2}-b^{2}=0$
Thus, only a data fission procedure performed with the true variance parameter (or at least an unbiased estimator of it) can ensure effective control of the Type I error.
Appendix 5 Supplementary Table 1
Conditional Fission Marginal Fission
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\right)$ $\sum\limits_{g=1}^{G}\pi_{g}\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)\left(\boldsymbol{\mu}_{g}-\sum\limits_{g=1}^{G}\pi_{g}\boldsymbol{\mu}_{g}\right)^{T}$ 0
$\operatorname{Cov}\left(\boldsymbol{X}_{i}^{(1)},\boldsymbol{X}_{i}^{(2)}\mid Z_{i}=g\right)$ 0 $\boldsymbol{\Sigma}_{g}-\boldsymbol{\Sigma}$
Table 1: Comparison of covariances under conditional and marginal data fission. The first row presents the covariance between the two fissioned random variables $\boldsymbol{X}_{i}^{(1)}$ and $\boldsymbol{X}_{i}^{(2)}$, while the second row shows the conditional covariance given class membership $Z_{i}=g$. Conditional fission preserves within-component independence, whereas marginal fission introduces bias due to the discrepancy between component-specific and global covariance structures.
Appendix 6 Supplementary Figure 1
Figure 1: Schematic view illustrating the circularity induced by conditional data fission for post-clustering differential analysis.